# Current-Enabled Optical Conductivity of Collective Modes in Unconventional Superconductors

### v2

@@ I Introduction
Determining the pairing symmetry and other properties of potentially unconventional superconductors remains a difficult task for many materials [1, 2, 3, 4, 5, 6]. One useful indicator of the nature of the superconducting ground state is the spectrum of collective modes. Although all superconductors can in principle host an amplitude and phase modes, usually referred to as the Higgs [7, 8, 9] and Nambu-Goldstone (NG) [10, 11] modes respectively, other collective modes can also be found in certain specific systems. One example is the Leggett mode [12], a relative phase oscillation between superconducting gaps, typically found in multiband superconductors. Another is the Bardasis-Schrieffer (BS) mode [13], which can occur when there are several competing superconducting pairing channels. The latter one corresponds to fluctuations of a subdominant order parameter. One issue with studying these collective modes is that they often do not couple linearly to light. A useful tool is Raman spectroscopy, which has been used to study the Higgs [14], Leggett [15], and BS modes [16]. The Higgs mode has also been detected using pump-probe spectroscopy [17, 18] and third harmonic generation [19, 20, 21].
All of these probes rely on nonlinear optical effects. In recent years, however, there has been several progress in predicting and observing signatures of collective modes of superconductors in the linear optical response. For example, the BS mode with a different parity than the ground state has been predicted to respond to light linearly in locally non-centrosymmetric compounds such as CeRh2As2 [22]. The Leggett mode has been predicted to become visible in the linear optical response in certain multiband superconductors [23, 24, 25]. Furthermore, the Higgs mode has been shown to become optically active in the presence of a supercurrent [26] (see also Refs. [27, 28]), which has been observed in an experiment [29]. This idea of injecting a supercurrent to induce a non-trivial optical response has also been considered in studying the effect of quasiparticle excitations in the optical conductivity [30, 31], where the supercurrent mainly plays a role of breaking inversion symmetry to allow the necessary optical transitions [31].
@@ II.1 One-band systems
$G^{-1}_{k,p}$ $=\mathcal{G}_{k}^{-1}-\Sigma^{e}_{k,p}-\Sigma^{\Delta}_{p},$ (11)
@@
Integrating out the fermions (see, for example, Ref. [38]) leaves us with an effective action, whose components that depend on the fluctuations of the order parameters and the vector potential $\bm{A}$ are given as
@@
where $n_{F}(E)$ is the Fermi-distribution. The terms in the effective action that depend on the vector potential $\bm{A}$ can be again found by performing a gaussian integral, which defines a new effective coupling-vertex between photons. Diagrammatically, this step is shown in Fig. 3, and one obtains the effective coupling,
@@ II.2 Multiband systems
where the off-diagonal elements of $V^{\mu}$ are responsible for mixing different bands, while the off-diagonal elements of $\Pi$ are responsible for mixing different pairing channels.
@@ III.1 One-band model
where $n_{F}^{\pm}=n_{F}(E^{+})-n_{F}(E^{-})$. The numerical solution for these equations are shown in Fig. 6. This plot is consistent with previous findings [32], where a continuum-model was used instead of a lattice. Calculations are always done at zero Temperature in the present work.
@@
In order to determine the full collective-mode spectrum, Eq. (26) has to be solved for $\Omega$. In Appendix B, this condition is simplified into more explicit mode equations for $\bm{q}=0$. The equations relevant to the BS Mode are given in Eqs. (90) and (92). The spectrum obtained from solving these equations numerically in the respective phase is shown in Fig. 7. In agreement with the previous findings [32], for such systems a BS mode is found in the $s$-wave phase, which softens at the transition to the $s+id$-wave ground state. In the $s+id$-wave phase, the mode consists of $s-id$-fluctuations, meaning out-of-phase amplitude fluctuations of the two order parameter components, which has been referred to as a mixed-symmetry Bardasis-Schrieffer (MSBS) mode [32], that stays below the gap throughout the phase. After the transition to a pure $d$-wave ground state the quasiparticle gap features nodes, and consequently there can be no collective excitations below the gap. [32]
@@ III.2 Two-band model
To explore multiband effects on the optical spectrum of the collective modes, we employ a simple extension of the one-band model used in the previous section. Instead of the two-dimensional single-layer model, we consider a bilayer model with a weaker single-particle interband hopping $J<t$. This corresponds to simply having two copies of the same cosine-dispersion band, placed at ±$\pm J-\mu$:
@@ III.2.1 The symmetric case
When $\mu=0$, the bands are centered at ±$\pm J$, and the multiband gap equation (See Appendix A) yields the same gap values in each band. The resulting quantum phase transition that takes place while tuning $V_{d}$ can be seen in Fig. 10. The small discontinuity at small $\Delta_{s}$ is a result of the gap equation being solved in the presence of a supercurrent.
@@ III.2.2 The asymmetric case
The situation changes qualitatively when the gaps for the two bands are different. Besides manually changing the interaction strength between the two bands, an easy way to achieve this is to set $\mu\neq 0$, which naturally leads to different gaps, $\Delta_{\mu-}$ and $\Delta_{\mu+}$. The phase diagram for $\mu=-0.2$ with the usual interband scattering [Eq. (31)] is shown in Fig. 14a.
Figure 14: Order-parameter components as a function of $|V_{d}|$ for the asymmetric bilayer model with the $s$- and $d$-wave pairing interactions as well as (a) with the interband pair scattering and (b) without the interband pair scattering [i.e., the off-diagonal elements of the matrix in Eq. (31) are set to zero]. The solid and dotted lines correspond to $\Delta_{s-}$ and $\Delta_{s+}$ (i.e., the gap in the lower and upper band), respectively. The parameters are $\mu=-0.2$, $V_{s}=-1$, $t=1$, and $\bm{q}=0.001\bm{\hat{e}}_{x}$.
The most notable difference to the symmetric case discussed in the last section is that the components of the order parameter now change discontinuously when entering the mixed symmetry phase. To understand the reason behind this, it is useful to consider the same phase diagram without any interband pair scattering, as shown in Fig. 14b. There, it can be seen that the two bands undergo the transition into the mixed symmetry state for different values of $V_{d}$. When Cooper pairs are allowed to scatter between the bands, the transitions in the two bands are forced to occur simultaneously. Because of this, when the $s+id$-wave state first becomes a solution of the self-consistence equation, the free energy of this solution is still above that of the pure $s$-wave state. Once the free energy of the $s+id$-wave state drops below that of the $s$-wave state, the ground state changes discontinuously to an $s+id$-state with a finite $id$-component.
@@
Figure 15: Collective-mode spectrum for the asymmetric bilayer model with the $s$- and $d$-wave pairing interactions. The four sub-gap modes are an out-of-phase BS/MSBS mode (green), a Leggett mode (red), an in-phase BS/MSBS mode (blue), and the NG mode (gray). The gaps in the lower (--) and upper (+)(+)( + ) bands are shown by the black solid and dotted lines, respectively. The parameters are $\mu=-0.2$, $V_{s}=-1$, $t=1$, and $\bm{q}=0.001\bm{\hat{e}}_{x}$.
he lower BS mode becomes gapless when the $s+id$-state first becomes a solution, but remains gapless until this state becomes the true ground state.
@@
Figure 16a shows how, as $|V_{d}|$ is varied, the peaks at each of the two gaps become much stronger and move to smaller frequencies, now being associated with the two BS modes rather than quasi-particle excitations. The peak due to the Leggett mode, which can only be seen in the inset of Fig. 16a, does not change. As the frequency of the lower BS mode approaches zero, the intensity of the corresponding peak becomes larger, while the other one becomes smaller. After the lower BS mode becomes gapless, the corresponding peak disappears from the optical spectrum, and only the upper BS mode and Leggett mode remain visible. After the transition to the $s+id$ ground state, the optical spectrum shown in Fig. 16b exhibits four peaks, for $V_{d}=-1.035$. The one at $\omega=0$ remains present throughout the $s+id$ phase as the Nambu-Goldstone mode, which appears to become optically active for the bilayer model in the time-reversal symmetry broken phase. In real situations, the NG mode is expected to be lifted to high energy by the Anderson-Higgs mechanism due to the coupling to electromagnetic fields [9], so this peak would not be observed.
The three other peaks, belonging to the symmetric MSBS, Leggett and asymmetric MSBS mode, respectively, are all expected to evade the Anderson-Higgs mechanism. As expected from Fig. 15, the upper two peaks eventually disappear from the sub-gap spectrum. This can be seen in Fig. 16b, for $V_{d}=1.105$. Initially, one only sees the peaks due to the lower MSBS and NG mode. Only in the inset can one see the very small peak at the lower quasiparticle gap, which remains as the strongly damped Leggett mode.
@@ IV.1.2 Superconductivity in Rashba systems
$H_{\text{int}}$ $=\sum_{\bm{kk^{\prime}}}\sum_{\{\lambda_{i}\}}t_{\lambda_{2}}(\bm{k})t_{\lambda_{3}}^{*}(\bm{k^{\prime}})\widetilde{V}^{\{\lambda_{i}\}}_{\bm{kk^{\prime}}}c^{\dagger}_{\bm{k}\lambda_{1}}c^{\dagger}_{\bm{-k}\lambda_{2}}c_{\bm{-k^{\prime}}\lambda_{3}}c_{\bm{k^{\prime}}\lambda_{4}},$ (44)
where the interaction is rescaled with the previously defined phase-factors (see Eq. (43)). The Hubbard-Stratonovich transformation is then performed exactly as before (see Eq. (3)) with the difference being an additional summation over the $\lambda$-indices. This gives
@@
The gap functions obey an equivalent self-consistency equation:
@@
Here $\tau^{-}_{\lambda_{3}\lambda_{4}}$ and ${G}$ are $4\times 4$ matrices written in the frequency-momentum representation as
@@
where $\widetilde{{G}}$ is given by ${G}$ with $\Delta$ being replaced with $\widetilde{\Delta}$. The entries of the gap function $\widetilde{\Delta}$ obey [35]
@@
Given that the following section will focus on the interband pairing, the important conclusions here are:
@@ IV.3 Results
The numerical solution of the gap equation in this system is shown in Fig. 17a for constant $V_{A_{1}}$ as a function of $V_{E}$, and in Fig. 17b for constant $V_{E}$ as a function of $V_{A_{1}}$. In each case, one can see a discontinuous transition between two phases, the nature of which can be identified from Eq. (57). The phase in which $V_{A_{1}}$ is dominant features a real $\bm{d}$-vector that lies in the plane. The phase in which $V_{E}$ is dominant has a time-reversal symmetry breaking, chiral order parameter of $p+ip$-type. Explicitly, the $\bm{d}$-vector takes the form of
@@
Figure 18: Collective-mode spectra for the Rashba model with interband pairing (a) in the real in-plane phase and (b) in the chiral out-of-plane phase. Each line corresponds to a collective mode. The $A_{1}$ phase (a) features only one NG mode (red) and one BS mode (blue). The chiral phase (b) features two BS modes: The red line corresponds to $A_{1}+iA_{1}$ fluctuations, and the blue one to $A_{1}-iA_{1}$ fluctuations. The other two modes correspond to relative phase oscillations $iE_{x}+E_{y}$ (purple) and out-of-phase amplitude oscillations $E_{x}-iE_{y}$ (green). Again, the parameters are $\mu=-3$, $t=1$, $\alpha=0.1$, $\bm{q}=0$, and $\omega_{c}=1$.
@@
The chiral phase (Fig. 18b) shows a more diverse spectrum: In addition to two non-degenerate BS modes (the red and blue lines) corresponding to fluctuations of the $A_{1}$ order parameter, it shows two more collective modes due to relative oscillations of the $E_{x}$ and $E_{y}$ order parameters, the frequency of which is independent of the strength of the sub-dominant in-plane order. We identify the lower one of these modes (green line in Fig. 18b) as the out-of-phase amplitude oscillations of the $p_{x}$ and $p_{y}$ components, while the upper mode corresponds to a relative phase oscillation between the components.
Given that this model breaks inversion symmetry, these collective modes are in principle allowed to contribute to the optical conductivity. Again, the particular system is chosen because it has been shown to feature Lifshitz invariants in the GL free energy [35], which are related to the optical response of collective modes [24]. In fact, this turned out to be the case for the BS modes, as seen in the numerical data in Fig. 19.
@@
For the chiral phase, the optical conductivity is shown in Fig. 20. As $V_{E}$ is increased, the BS modes found in Fig. 18b generate peaks in the optical conductivity. The most notable difference with respect to all the other optical responses studied so far is that here, the modes also result in a nonzero transverse optical conductivity (i.e., the ac Hall conductivity). Satisfying $\sigma_{xy}=-\sigma_{yx}$, these parts of the optical conductivity have opposite signs on each side of the mode peak. Exactly at the mode frequency, the components vanish. Such peaks can be measured in experiments via the polar Kerr effect, where the polarization angle of a reflected ray of light becomes different from that of the incident light [46, 47].

### v3

@@ I Introduction
Determining the pairing symmetry and other properties of potentially unconventional superconductors remains a difficult task for many materials [1, 2, 3, 4, 5, 6]. One useful indicator of the nature of the superconducting ground state is the spectrum of collective modes. Although every superconductor can in principle host an amplitude and phase modes, usually referred to as the Higgs [7, 8, 9] and Nambu-Goldstone (NG) [10, 11] modes respectively, other collective modes can also be found in certain specific systems. One example is the Leggett mode [12], a relative phase oscillation between superconducting gaps, typically found in multiband superconductors  [^1] In the present work, we are interested in spatially uniform oscillations of the relative phase of the superconducting order parameters. There are, however, also situations where a stable state with a topological winding of the relative phase in space can exist, which is called a phase soliton [53, 54, 55].. Another collective excitation can occur when there are several competing superconducting pairing channels. In those systems, where the ground state belongs to one pairing channel but other attractive channels exist, it is possible to find well-defined excitations below the quasi-particle gap. These can be identified as fluctuations of the order parameter belonging to the sub-dominant pairing channel, and are referred to as the Bardasis-Schrieffer (BS) modes [14].
One issue with studying these collective modes is that they often do not couple linearly to light. A useful tool is Raman spectroscopy, which has been used to study the Higgs [15], Leggett [16], and BS modes [17]. The Higgs mode has also been detected using pump-probe spectroscopy [18, 19] and third harmonic generation [20, 21, 22]. All of these probes rely on nonlinear optical effects. In recent years, however, there has been several progress in predicting and observing signatures of collective modes of superconductors in the linear optical response. For example, the BS mode with a different parity than the ground state has been predicted to respond to light linearly in locally non-centrosymmetric compounds such as CeRh2As2 [23]. The Leggett mode has been predicted to become visible in the linear optical response in certain multiband superconductors [24, 25, 26]. Furthermore, the Higgs mode has been shown to become optically active in the presence of a supercurrent [27] (see also Refs. [28, 29]), which has been observed in an experiment [30]. This idea of injecting a supercurrent to induce a non-trivial optical response has also been considered in studying the effect of quasiparticle excitations in the optical conductivity [31, 32], where the supercurrent mainly plays a role of breaking inversion symmetry to allow the necessary optical transitions [32].
@@ II.1 One-band systems
$G^{-1}_{k,p}$ $=\mathcal{G}_{k}^{-1}-\Sigma^{e}_{k,p}-\Sigma^{\Delta}_{k,p},$ (11)
@@
Integrating out the fermions (see, for example, Ref. [39]) leaves us with an effective action, whose components that depend on the fluctuations of the order parameters and the vector potential $\bm{A}$ are, with $\Sigma=\Sigma^{e}+\Sigma^{\Delta}$:
@@
where $n_{F}(E)$ is the Fermi-Dirac-distribution. The terms in the effective action that depend on the vector potential $\bm{A}$ can be again found by performing a gaussian integral, which defines a new effective coupling-vertex between photons. Diagrammatically, this step is shown in Fig. 3, and one obtains the effective coupling,
@@ II.2 Multiband systems
where the off-diagonal elements of $V^{\mu}$ are responsible for mixing different bands, while the off-diagonal elements of $\Pi$ are responsible for mixing different pairing channels. The same condition Eq. 26 applies here as in the single-band case.
@@ III.1 One-band model
where $n_{F}^{\pm}=n_{F}(E^{+})-n_{F}(E^{-})$. The numerical solution for these equations are shown in Fig. 6. This plot is consistent with previous findings [33], where a continuum-model was used instead of a lattice. Calculations are always done at zero Temperature in the present work. To confirm the fact that the mixed symmetry state becomes the true ground state, we checked that it has a lower free energy than the pure $s$-wave ground state, see Appendix E.
@@
In order to determine the full collective-mode spectrum, Eq. (26) has to be solved for $\Omega$. In Appendix B, this condition is simplified into more explicit mode equations for $\bm{q}=0$. The equations relevant to the BS Mode are given in Eqs. (94) and (96). The spectrum obtained from solving these equations numerically in the respective phase is shown in Fig. 7. In agreement with the previous findings [33], for such systems a BS mode is found in the $s$-wave phase, which softens at the transition to the $s+id$-wave ground state. In the $s+id$-wave phase, the mode consists of sids-id-fluctuations, meaning out-of-phase amplitude fluctuations of the two order parameter components, which has been referred to as a mixed-symmetry Bardasis-Schrieffer (MSBS) mode [33], that stays below the gap throughout the phase. After the transition to a pure $d$-wave ground state the quasiparticle gap features nodes, and consequently there can be no collective excitations below the gap. [33]
@@ III.2 Two-band model
To explore multiband effects on the optical spectrum of the collective modes, we employ a simple extension of the one-band model used in the previous section. Instead of the two-dimensional single-layer model, we consider a bilayer model with a weaker single-particle interband hopping $J<t$. This corresponds to simply having two copies of the same cosine-dispersion band, placed at $\pm J-\mu$:
@@ III.2.1 The symmetric case
When $\mu=0$, the bands are centered at $\pm J$, and the multiband gap equation (See Appendix A) yields the same gap values in each band. The resulting quantum phase transition that takes place while tuning $V_{d}$ can be seen in Fig. 10. The small discontinuity at small $\Delta_{s}$ is a result of the gap equation being solved in the presence of a supercurrent. Again, we confirmed the transition with a free energy comparison, see Appendix E.
@@ III.2.2 The asymmetric case
The situation changes qualitatively when the gaps for the two bands are different. Besides manually changing the interaction strength between the two bands, an easy way to achieve this is to set $\mu\neq 0$, which naturally leads to different gaps, $\Delta_{\mu-}$ and $\Delta_{\mu+}$. The phase diagram for $\mu=-0.2$ with the usual interband scattering [Eq. (31)] is shown in Fig. 14.
Figure 14: Order-parameter components as a function of $|V_{d}|$ for the asymmetric bilayer model with the $s$- and $d$-wave pairing interactions. The solid and dotted lines correspond to $\Delta_{s-}$ and $\Delta_{s+}$ (i.e., the gap in the lower and upper band), respectively. The parameters are $\mu=-0.2$, $V_{s}=-1$, $t=1$, and $\bm{q}=0.001\bm{\hat{e}}_{x}$.
The most notable difference to the symmetric case discussed in the last section is that the order parameter now takes different values in the two bands.
@@
Figure 15: Collective-mode spectrum for the asymmetric bilayer model with the $s$- and $d$-wave pairing interactions. The four sub-gap modes are an out-of-phase BS/MSBS mode (green), a Leggett mode (red), an in-phase BS/MSBS mode (blue), and the NG mode (gray). The gaps in the lower (-) and upper $(+)$ bands are shown by the black solid and dotted lines, respectively. The parameters are $\mu=-0.2$, $V_{s}=-1$, $t=1$, and $\bm{q}=0.001\bm{\hat{e}}_{x}$.
The lower BS mode again becomes gapless at the transition point to the $s+id$ ground state.
@@
Figure 16a shows how, as $|V_{d}|$ is varied, the peaks at each of the two gaps become much stronger and move to smaller frequencies, now being associated with the two BS modes rather than quasi-particle excitations. The peak due to the Leggett mode, which can only be seen in the inset of Fig. 16a, does not change. As the frequency of the lower BS mode approaches zero, the intensity of the corresponding peak becomes larger, while the other one becomes smaller. After the transition to the $s+id$ ground state, the optical spectrum shown in Fig. 16b exhibits three peaks, for $V_{d}=-1.035$. As expected from Fig. 15, the upper two peaks, corresponding to the out of phase MSBS and Leggett mode, eventually disappear from the sub-gap spectrum. This can be seen in Fig. 16b, for $V_{d}=1.105$. Here, only a broad peak above the gap remains of these dampened collective modes, which is visible only in the inset.
@@ IV.1.2 Superconductivity in Rashba systems
$H_{\text{int}}$ $=\sum_{\bm{kk^{\prime}}}\sum_{\{\lambda_{i}\}}t_{\lambda_{2}}(\bm{k})t_{\lambda_{3}}^{*}(\bm{k^{\prime}})\widetilde{V}^{\{\lambda_{i}\}}_{\bm{kk^{\prime}}}a^{\dagger}_{\bm{k}\lambda_{1}}a^{\dagger}_{\bm{-k}\lambda_{2}}a_{\bm{-k^{\prime}}\lambda_{3}}a_{\bm{k^{\prime}}\lambda_{4}},$ (44)
where the interaction is rescaled with the previously defined phase-factors (see Eq. (43)). The Hubbard-Stratonovich transformation is then performed exactly as before (see Eq. (3)) with the difference being an additional summation over the $\lambda$-indices. Note that the fermionic fields will correspond to the $a$-operators, meaning they are expressed in the diagonal basis of the non-interacting hamiltonian. This gives
@@
Note that when we speak of interband-pairing, this corresponds to the off-diagonal elements of this matrix, connecting the ++ and the - band with each other. Since the Fermi surfaces of these bands are different, this would likely lead to finite-momentum cooper pairs in the groundstate of a physical system. For our purposes, to separate the effects of the cooper pair momentum and the inversion-breaking of the SOC, we assume that the cooper pairs have zero momentum in the ground state before applying a supercurrent. The gap functions obey an equivalent self-consistency equation:
@@
Here $\tau^{-}_{\lambda_{3}\lambda_{4}}$ and G{G} are $4\times 4$ matrices written in the frequency-momentum representation as
@@
where $\widetilde{{G}}$ is given by G{G} with $\Delta$ being replaced with $\widetilde{\Delta}$. The entries of the gap function $\widetilde{\Delta}$ obey [36]
@@
Here, we have renamed $\widetilde{\Delta}^{\text{inter}}_{\bm{k}}\equiv\widetilde{\Delta}^{+-}_{\bm{k}}=-\widetilde{\Delta}^{-+}_{\bm{-k}}$. Given that the following section will focus on the interband pairing, the important conclusions here are:
@@ IV.3 Results
The numerical solution of the gap equation in this system is shown in Fig. 17a for constant $V_{A_{1}}$ as a function of $V_{E}$, and in Fig. 17b for constant $V_{E}$ as a function of $V_{A_{1}}$. In each case, one can see a discontinuous transition between two phases, the nature of which can be identified from Eq. (57). The transition point was determined by comparing the free energy of the different solutions to the self-consistence equation, see Appendix E. The phase in which $V_{A_{1}}$ is dominant features a real $\bm{d}$-vector that lies in the plane. The phase in which $V_{E}$ is dominant has a time-reversal symmetry breaking, chiral order parameter of $p+ip$-type. Explicitly, the $\bm{d}$-vector takes the form of
@@
Figure 18: Collective-mode spectra for the Rashba model with interband pairing (a) in the real in-plane phase and (b) in the chiral out-of-plane phase. Each line corresponds to a collective mode. The $A_{1}$ phase (a) features only one NG mode (red) and one BS mode (blue). The chiral phase (b) features two BS modes: The red line corresponds to a ”left-winding” BS-mode of $e^{+i\omega t}A_{1}$-fluctuations, and the blue one to a ”right-winding” mode with $e^{-i\omega t}A_{1}$-fluctuations. The other two modes in purple and green correspond to linear superpositions of relative phase oscillations and relative amplitude oscillations. Again, the parameters are $\mu=-3$, $t=1$, $\alpha=0.1$, $\bm{q}=0$, and $\omega_{c}=1$.
@@
The chiral phase (Fig. 18b) shows a more diverse spectrum: In addition to two non-degenerate BS modes (the red and blue lines) corresponding to fluctuations of the one-component $A_{1}$ order parameter, it shows two more collective modes due to relative oscillations of the $E_{x}$ and $E_{y}$ order parameters, the frequency of which is independent of the strength of the sub-dominant in-plane order. From the eigenvectors of the effective coupling matrix at the mode frequency, we can identify different fluctuating components and their relative phase shifts. This yields the following expressions for the 4 differently colored modes in Fig. 18b, expressing $\delta\Delta(t)$ in real time in terms of $\widehat{\varphi}_{A_{1}}$, $\widehat{\varphi}_{E_{x}}$ and $\widehat{\varphi}_{E_{y}}$:
$\text{Red: }\delta\Delta$ $\propto\widehat{\varphi}_{A_{1}}e^{+i\omega t}$ (71)
$\text{Blue: }\delta\Delta$ $\propto\widehat{\varphi}_{A_{1}}e^{-i\omega t}$ (72)
$\begin{split}\text{Purple: }\delta\Delta&\propto 2\left[\widehat{\varphi}_{E_{x}}-i\widehat{\varphi}_{E_{y}}\right]\cos(\omega t)\\ &\quad-\left[\widehat{\varphi}_{E_{x}}+i\widehat{\varphi}_{E_{y}}\right]\sin(\omega t)\end{split}$ (73)
$\begin{split}\text{Green: }\delta\Delta&\propto\left[\widehat{\varphi}_{E_{x}}-i\widehat{\varphi}_{E_{y}}\right]\cos(\omega t)\\ &\quad+2\left[\widehat{\varphi}_{E_{x}}+i\widehat{\varphi}_{E_{y}}\right]\sin(\omega t)\end{split},$ (74)
where $\omega$ is the mode frequency which is of course different for each mode. Given that this model breaks inversion symmetry, these collective modes are in principle allowed to contribute to the optical conductivity. Again, the particular system is chosen because it has been shown to feature Lifshitz invariants in the GL free energy [36], which are related to the optical response of collective modes [25]. In fact, this turned out to be the case for the BS modes, as seen in the numerical data in Fig. 19.
@@
For the chiral phase, the optical conductivity is shown in Fig. 20. As $V_{A_{1}}$ is increased, the BS modes found in Fig. 18b generate peaks in the optical conductivity. The most notable difference with respect to all the other optical responses studied so far is that here, the modes also result in a nonzero transverse optical conductivity (i.e., the ac Hall conductivity). Satisfying $\sigma_{xy}=-\sigma_{yx}$, these parts of the optical conductivity have opposite signs on each side of the mode peak. Exactly at the mode frequency, the components vanish. Such peaks can be measured in experiments via the polar Kerr effect, where the polarization angle of a reflected ray of light becomes different from that of the incident light [47, 48].